OpenStudy (anonymous):

Find the range of the function: (3)/(sq.rt 9-x^2)

myininaya (myininaya):

\[y=\frac{3}{\sqrt{9-x^2}}\] Is that correct?

OpenStudy (anonymous):

yes

myininaya (myininaya):

\[y \sqrt{9-x^2}=3 \text{ note: multiply } \sqrt{9-x^2} \text{ on both sides}\] \[y^2(9-x^2)=9 \text{ note: squared both sides}\] \[9y^2-y^2x^2=9 \text{ note: distribute on left hand side}\] \[-y^2x^2=9-9y^2 \text{ note: subtracted } 9y^2 \text{ on both sides }\] \[x^2=\frac{9-9y^2}{-y^2} \text{ note: divided } -y^2 \text{ on both sides}\] \[x=\pm \sqrt{\frac{9-9y^2}{-y^2}} \text{ take square root of both sides }\] So tell me what y can be so that x exist? y cannot be 0 since we would have 0 in the bottom if y could be 0 we have that negative in the bottom I'm going to rewrite x a little more... \[x=\pm \sqrt{\frac{-(9-9y^2)}{y^2}}=\pm \sqrt{\frac{9y^2-9}{y^2}}\] So y^2 is already positive we want the inside to be either 0 or positive In other words we want \[9y^2-9 \ge 0\] \[9(y^2-1) \ge 0\] \[y^2-1 \ge 0\] \[(y-1)(y+1) \ge 0\] So (y-1)(y+1)=0 when y=-1 or y=1 But remember y cannot be negative look at the function we started out with both top and bottom are both positive for all x |------|------------ 0 1 We need to test these two intervals We have two intervals to test The first one is (0,1) The second is (1,inf) Interval 1: Choose some value in that interval evaluate (y-1)(y+1) Interval 2 : Choose some value in that interval Do the same as before plug into (y-1)(y+1)

myininaya (myininaya):

So I solved for x to see what y could be

OpenStudy (anonymous):

thanks so much myininaya